OpenStudy (anonymous):

Worked this on my own, not sure of my answer is correct....? What is a possible first step to eliminate the variable y in the following system?4x + 6y = 15x - 3y = 8 Multiply the first equation by 5. Multiply the first equation by -2. Multiply the second equation by 2. Multiply the second equation by -3.

OpenStudy (anonymous):

multiply the second equation by 2

OpenStudy (anonymous):

I got b, though? why not -2?

OpenStudy (anonymous):

then, you can subtract both the equations easily eliminating the y-term

OpenStudy (anonymous):

Sorry, add

OpenStudy (anonymous):

If you do as in b, then you will have to subtract the equations and you will have one extra step than doing as in c.

OpenStudy (anonymous):

Hey, wait...i thought "b" to multiply second equation(didn't have a proper look) by -2.

OpenStudy (anonymous):

I am sorry...now i'll show you what happens if you proceed as in b...

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

If you multiply the first equation by -2, it becomes -8x-12y= -16 and we have second equation as 15x-3y=8. its not possible to solve it unless you multiply second equation by 4. Its quite lengthy...

OpenStudy (anonymous):

Oh, i see what you're saying

OpenStudy (anonymous):

So the question is asking for the most ACCURATE, yet SIMPLE approach?

OpenStudy (anonymous):

If you do as in c, the second equation becomes 30x-6y=16 and the first equation as given is 4x+3y=8 and now its easy to solve...yes simple and accurate approach

OpenStudy (anonymous):

4x+6y=8 i made a typo above..

OpenStudy (anonymous):

The answer I get is C

OpenStudy (anonymous):

4x+6y=8 is not one of the equations.

OpenStudy (anonymous):

So, you are good..you are the first person on this site who wanna learn, as far as i've experienced...yes its C

OpenStudy (anonymous):

awwww thank you!!!!!!:))))

OpenStudy (anonymous):

@CliffSedge According to your master theory, you are right but I gave a solution to the question represented irrespective of the error made by her professor

OpenStudy (anonymous):

Which is? @AbhimanyuPudi

OpenStudy (anonymous):

Yes. Your answer was correct, but I wanted to make sure we all know what's really going on here.

OpenStudy (anonymous):

well, the error as found out by our partner is that they are not compound equations..